OpenStudy (anonymous):

The width of a rectangle is one-third the length. If the perimeter of the rectangle is 56 units, what is the area in square units?

OpenStudy (math&ing001):

I feel nostalgic to these kind of problems, they were fun =) So width =1/3 length and perimeter = 56 You get a system : 2*( width + length) = 56 and width =1/3 length Let's plug in width from 1st equation by its value from 2nd equation, what do you get ?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

I don't understand

OpenStudy (math&ing001):

Ok what part is confusing you ? Let x be the width and y be the length. The width of a rectangle is one-third the length. So x=1/3 y Up till here good ?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so is it 2(1/3 + l) =56, like that

OpenStudy (math&ing001):

Yeah, but don't forgot about the variable : 2*( x + y ) = 56 2*( 1/3 y + y) =56 2*( 1/3 + 1) y = 56 From here you just solve for y.

OpenStudy (anonymous):

ok thanks